AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
DESPEC
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Thu Apr 3 19:29:09 2025
Author
*
:
Subject
*
:
08:15 beam tuning in progress 08.30 System wide checks - aida06 global clock failure (check clock status) - aida02 , aida04, aida06 fail calibration (ADC calibration) - SYNC error counter ok - SYNC pulse check ok - 'collect memory information from FEE64 Linux' still returns scripting error -> will not check again until update - WR decoder test passes - good event statistics OK, except aida 09, 10, 11, 12 as yesterday morning (see attach 1) - FEE temperature OK (see attach 2) - Detector OK (see attach 3) - NewMerger. MergerHTTPd, newMerger Terminal, data relay MBS screenshots attached (4,5,6,7) 10:37 Issues with magnets due to the new control system. Beam tuning continues AIDA continues running OK - aida06 global clock failure (check clock status) - aida02 , aida04, aida06 fail calibration (ADC calibration) - SYNC error counter ok - SYNC pulse check ok - WR decoder test passes - good event statistics OK, except aida 09, 10 (see attach 8). Aida 9 & 12 rates now lower - FEE temperature OK - Detector OK - Merger OK 12:21 Beam start AIDA continues running OK - aida06 global clock failure (check clock status) - aida02 , aida04, aida06 fail calibration (ADC calibration) - SYNC error counter ok - SYNC pulse check ok - WR decoder test passes -ASIC check load performed -Writing to file R11 in directory 290319 -histograms zeroed -good event statistics OK, except aida 09, 10 (attachment 9) -FEE temperatures ok (attachment 10) -detector bias ok (attachment 11) -thick degrader was placed in front of snout during beam operation 12:40 beam stop TapeServer stopped, file R11_0 in 290319 TapeServer restarted, not writing to file low energy spectra for aida09 (attachment 12) high energy spectra for aida09 (attachment 13) high energy spectra for aida11 (attachment 14) 1D hit patterns (stat) in aida09,10,11,12 (attachment 15) ASIC control settings used during beam, including thresholds (attachment 16) 14.20 *NOTE degrader setting different from yesterday, silver beam still not observed, just unidentified light ions. AIDA continues running OK aida06 global clock failure (check clock status) aida02 , aida04, aida06 fail calibration (ADC calibration) SYNC error counter ok SYNC pulse check ok WR decoder test passes 14.37 good event statistics OK, except aida 10 (attachment 17) detector bias ok (attachment 18) FEE temperatures ok (attachment 19) 14.54 Degrader removed 16:13 AIDA continues running OK aida06 global clock failure (check clock status) aida02 , aida04, aida06 fail calibration (ADC calibration) SYNC error counter ok SYNC pulse check ok WR decoder test passes FEE64 temperatures OK good event statistics OK, except aida 3, 9, 10 running hot (see attach 18) 17:00 Beam start - no degrader AIDA running OK Writing on file R13 Beam is better focused (see attach 19). Appears to be off-centre in the bottom right part of the detector Rates OK (see attach 20) DSSD bias/current OK (see attach 21) 17:21 Beam re-centred. Histograms cleared - see attach 22 AIDA continues OK Run continues High energy spectra in aida09 (attachment 25) Beam rate approx 5000 ions/s Beam being diffused TapeServer start stop - AIDA writing to file R14 in 290319 17:48 1D hit patterns - beam in process of being diffused (attachment 26) 17:57 leakage current back at 1.465 uA (attachment 27) 18:08 beam stopped TapeServer stopped and restarted - not writing to file updated hit patterns from beam diffusing (attachment 28) aida10, ADC3 high energy spectra (attachment 29) 18.11 calibrating tpc with masks TapeServer stop/start R16, 290319 18.49 leakage current briefly fluctuating between ~1.2 uA and ~2 uA, but has returned to stability at ~ 1.4 uA (attachment 30) 18.55 beam stopped for tuning - TapeServer stop/start. Not writing to file. 18.58 leakage current still stable after beam off (attachment 31) 19.10 beam start. TapeServer stop/start writing to file R18 in 290319 19.15 leakage current still ok after beam activation 19.34 fee12 high energy spectra (attachment 32) 20.35 beam stop, tapeserver stopped histogramming link dropped on aida 12 20.38 beam start. TapeServer start R20
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown